# Ergotropy transport in a one dimensional spin chain

### v1

@@ III.3 Robustness to disorder
While the previous sections considered non-PST couplings, with a particular focus on uniform couplings, we now focus on PST. We will once again compare the ability for the chain to shuttle ergotropy depending on if the how initial extractable work is encoded. We will consider the PST coupling regime ($\alpha=1$) and disorder in the coupling strength $\mathcal{J}_{\alpha}$. The disorder is drawn from a uniform distribution in the range $\Delta,+\Delta]$ Ronke et al. (2011); Alsulami et al. (2022); Kiely and Campbell (2021); Coopmans et al. (2022), where $\Delta$ quantifies the disorder strength.
@@
In Fig. 4(a) we show $\langle\varepsilon_{max}\rangle$ obtained at the first reflection point where $\langle\cdot\rangle$ denotes the ensemble average over $1000$ disorder realizations. We can see clearly see that for all system sizes and disorder strengths, the coherent superposition outperforms the mixed state. The superposition case also exhibits smaller deviations from the mean value, as captured by the size of the shaded regions which correspond to one standard deviation. We remark that the initial quadratic drop in ergotropy can be understood and quantified using either master equations for the ensemble average state Kropf et al. (2016); Kiely (2021) or the use of perturbation theory Poggi et al. (2024).
@@ IV Thermodynamic Cost of Quench Protocol
In the preceding sections, we have examined the transport of extractable work, which as we have shown has subtle differences with the transport of population. However, there is a work cost associated with the sudden quench of the interactions needed for this transport to occur. In the next section, we will quantify this thermodynamic cost using the well known quantum work statistics Talkner et al. (2007); Silva (2008); Santini et al. (2023); Zawadzki et al. (2020, 2023); Kiely et al. (2023); Apollaro et al. (2015).
We assume an initial Hamiltonian $\mathcal{H}^{i}$ where the couplings are switched off (i.e. Eq. (2) with $\mathcal{J}_{\alpha}\!=\!0$) and a final Hamiltonian $\mathcal{H}^{f}$ with the couplings on. Assuming the spectral decomposition for these Hamiltonians, i.e. $\mathcal{H}^{i,f}\!=\!\sum_{m}E_{m}^{i,f}|E_{m}^{i,f}\rangle\langle E_{m}^{i,f}|$, we can evaluate work involved in suddenly quenching on the interaction via the two point measurement (TPM) protocol Talkner et al. (2007). The TPM protocol provides us access to the quantum work distribution,
@@
For a given initial state, $\varrho$, the first measurement has probability $p_{m}\!=\!\langle E_{m}^{i}|\varrho|E_{m}^{i}\rangle$. The conditional probability for the second measurement follows as $p_{k|m}\!=\!\langle E_{k}^{f}|\varrho|E_{m}^{i}\rangle$, where we have already assumed the sudden quench limit where the interactions are switched on effectively instantaneously. The observed values of work correspond to energy differences from the two measurements, $W_{k,m}\!=\!E_{k}^{f}-E_{m}^{i}$. For the initial states considered in this work, Eqs. (9) and (10), it follows that for $p_{m}$ there are at most only two possible non-zero values. From Eq. (19) the moments of the work distribution follow as $\langle W^{n}\rangle\!=\!\int dWW^{n}P(W)$, which are a commonly used summary statistic Kiely et al. (2023); Zawadzki et al. (2023).
@@
It is worth commenting that the preceding analysis also applies more generally to assessing the thermodynamics of quantum communication protocols De Chiara et al. (2005); Ronke et al. (2011); Kay (2010). Our results demonstrate that at the level of the work done in switching on the necessary interactions, there are notable differences that arise depending on the precise distribution of couplings employed. Interestingly, we see that ideal communication channels exhibit smaller fluctuations in the work cost of turning on the couplings.
@@ Appendix A XX PST Coupling Analytics
Here we work in the single-excitation subspace, following the results of Albanese et al. (2004). This allows us to obtain analytical expressions for the transition amplitude function and the work distribution when the XX model is tuned with PST coupling. The nearest-neighbor hopping Hamiltonian for a chain with position-dependent couplings is given by
@@
with the natural boundary conditions $\psi(0)=\psi(N+1)=0.$ This recurrence relation can be shown to be equivalent to an eigenvalue equation for Krawtchouk polynomials Christandl et al. (2017),
@@ B.2 Work distribution
for $2J\leq W\leq 2J$ and $p(W,J)=0$ otherwise. While previous works have studied work statistics in an XX spin chain, for example Moriya et al. (2019), the analytic expression for the probability density function in Eq.(53) has not been reported. This distribution has a zero first moment of work and a second moment $\langle W^{2}\rangle=J^{2}$.

### v2

@@ III.3 Robustness to disorder
While the previous sections considered non-PST couplings, with a particular focus on uniform couplings, we now focus on PST. We will once again compare the ability for the chain to shuttle ergotropy depending on if the how initial extractable work is encoded. We will consider the PST coupling regime ($\alpha=1$) and disorder in the coupling strength $\mathcal{J}_{\alpha}$. The disorder is drawn from a uniform distribution in the range $\Delta,+\Delta]$ Ronke2011PRA; IreneDisorder2; magnonTransport; Coopmans2022, where $\Delta$ quantifies the disorder strength. We remark that for weak disorder strengths, the impact of the noise is dictated by the variance, rather than the particular choice, of the distribution DisorderPRX. Therefore, we choose a uniform distribution as it effectively represents a worst case scenario as the strength of the disorder is increased IreneDisorder2.
@@
In Fig. 4(a) we show $\langle\varepsilon_{max}\rangle$ obtained at the first reflection point where $\langle\cdot\rangle$ denotes the ensemble average over $1000$ disorder realizations. We can see clearly see that for all system sizes and disorder strengths, the coherent superposition outperforms the mixed state. The superposition case also exhibits smaller deviations from the mean value, as captured by the size of the shaded regions which correspond to one standard deviation. We remark that the initial quadratic drop in ergotropy can be understood and quantified using either master equations for the ensemble average state DisorderPRX; kiely2021 or the use of perturbation theory poggi2024.
@@ IV Thermodynamic Cost of Quench Protocol
In the preceding sections, we have examined the transport of extractable work, which as we have shown has subtle differences with the transport of population. However, there is a work cost associated with the sudden quench of the interactions needed for this transport to occur. In the next section, we will quantify this thermodynamic cost using the well known quantum work statistics WorkNotObservable; Silva2008; Santini2023; Zawadzki2020; NonGaussworkStat; EntQWD; apollaro2015work.
We assume an initial Hamiltonian $\mathcal{H}^{i}$ where the couplings are switched off (i.e. Eq. (2) with $\mathcal{J}_{\alpha}\!=\!0$) and a final Hamiltonian $\mathcal{H}^{f}$ with the couplings on. Assuming the spectral decomposition for these Hamiltonians, i.e. $\mathcal{H}^{i,f}\!=\!\sum_{m}E_{m}^{i,f}|E_{m}^{i,f}\rangle\langle E_{m}^{i,f}|$, we can evaluate work involved in suddenly quenching on the interaction via the two point measurement (TPM) protocol WorkNotObservable. The TPM protocol provides us access to the quantum work distribution,
@@
For a given initial state, $\varrho$, the first measurement has probability $p_{m}\!=\!\langle E_{m}^{i}|\varrho|E_{m}^{i}\rangle$. The conditional probability for the second measurement follows as $p_{k|m}\!=\!\langle E_{k}^{f}|\varrho|E_{m}^{i}\rangle$, where we have already assumed the sudden quench limit where the interactions are switched on effectively instantaneously. The observed values of work correspond to energy differences from the two measurements, $W_{k,m}\!=\!E_{k}^{f}-E_{m}^{i}$. For the initial states considered in this work, Eqs. (9) and (10), it follows that for $p_{m}$ there are at most only two possible non-zero values. From Eq. (19) the moments of the work distribution follow as $\langle W^{n}\rangle\!=\!\int dWW^{n}P(W)$, which are a commonly used summary statistic EntQWD; NonGaussworkStat.
@@
It is worth commenting that the preceding analysis also applies more generally to assessing the thermodynamics of quantum communication protocols deChiara2005PRA; Ronke2011PRA; kay2010perfect. Our results demonstrate that at the level of the work done in switching on the necessary interactions, there are notable differences that arise depending on the precise distribution of couplings employed. Interestingly, we see that ideal communication channels exhibit smaller fluctuations in the work cost of turning on the couplings.
@@ Appendix A XX PST Coupling Analytics
Here we work in the single-excitation subspace, following the results of Albanese2004. This allows us to obtain analytical expressions for the transition amplitude function and the work distribution when the XX model is tuned with PST coupling. The nearest-neighbor hopping Hamiltonian for a chain with position-dependent couplings is given by
@@
with the natural boundary conditions $\psi(0)=\psi(N+1)=0.$ This recurrence relation can be shown to be equivalent to an eigenvalue equation for Krawtchouk polynomials Christandl2017,
@@ B.2 Work distribution
for $2J\leq W\leq 2J$ and $p(W,J)=0$ otherwise. While previous works have studied work statistics in an XX spin chain, for example Moriya2019, the analytic expression for the probability density function in Eq.(53) has not been reported. This distribution has a zero first moment of work and a second moment $\langle W^{2}\rangle=J^{2}$.